# Statistics of topological defects across a phase transition in a superconducting quantum processor

### v1

@@ A.1 Noise renormalization
The workhorse of our error mitigation protocol is based on a noise renormalization technique, successfully used in [43, 44, 22, 24, 45]. The idea is to assume a noise model, estimate the corresponding parameters on the quantum device and reverse its effect in the post-processing stage. In the following, we shall assume a depolarising noise model $\mathcal{N}_{p}[\cdot]$ with parameter $p$, which maps the expectation value of a Pauli observable $\sigma$ under a quantum state $\rho$ to
@@
If the parameter $p$ is known, the results can be corrected by dividing the noisy results by $(1-p)$. The main idea is then to estimate $(1-p)$ by running a circuit with known expectation value. This becomes particularly simple when $h=0$ or $th=\pi$. Hence in the first case the evolution of the field is the identity and in the second a uniform bit flip. In both cases, this action commutes with the interaction part, and we have
+¯$\langle\bar{+}|\mathcal{U}^{\dagger}(t)\hat{n}\mathcal{U}(t)|\bar{+}\rangle=1,$ (12)
where $|\bar{+}\rangle$ is the uniform superposition state obtained by applying a layer of Hadamard gates to $|0\rangle^{\otimes N}$. Since we require that the circuit is as close as the original one as possible, we choose the field at time $t$ such as $h=\pi/t$. Even if the expectation values takes a simple form, the full circuit is still executed, enabling us to effectively estimate $(1-p)$ by evaluating Eq. 12. We note that, in practice, each component of $\hat{n}$ are individually re-normalized.
A.2 Randomized compiling
[…]
Twirling can also be used to mitigate measurement errors [47] in the computation of expectation values. This process is performed by randomly flipping the state of the qubit before measurement, and reverse the value of the measured bit, if applicable. effectively, measurement twirling diagonalise the readout transfer-matrix, making it easier to invert using techniques describe below.
A.3 Readout error mitigation
We mitigate readout errors by calibrating the device. Since a full calibration is exponentially expensive, we adopt a sparse strategy [48, 61] by only measuring the $N$-qubit state $|0\rangle^{\otimes N}$ and $|1\rangle^{\otimes N}$, and build the confusion matrices
@@ A.3 Readout error mitigation
As stated above, measurement twirling is performed to diagonalize the transfer-matrix, thus making it easier to invert.
Appendix B Propagation of statistical uncertainties
@@
Here, $\alpha$, $\beta>0$ parameterize the beta distribution and represent the number of times the basis vector $i$ is measured. Note that we only sample from binary string appearing at least once in the data, to avoid sampling from exponentially many terms. After this inference phase, the following procedure can be used to determine the expected values: (i) sample a value $p_{k}^{\prime}$ from the posterior $P(p_{k}^{\prime}|m_{i})$. (ii) sample $L$ new measurements from the likelihood $P_{b}(m_{k}^{\prime};p_{k}^{\prime})$. (iii) Compute expectations values by averaging over the generate measurements. To generalized to multiple qubits, we can use the Dirichlet distribution, which is a closed-form prior for a multinominal distribution.

### v2

@@ II.3 Error mitigation
The workhorse of our error mitigation protocol is based on a noise renormalization technique, successfully used in Urbanek et al. (2021); A Rahman et al. (2022); Farrell et al. (2024); Kiss et al. (2025a); Kavaki and Lewis (2024). The idea is to assume a noise model, estimate the corresponding parameters on the quantum device and reverse its effect in the post-processing stage. In the following, we shall assume a depolarising noise model $\mathcal{N}_{p}[\cdot]$ with parameter $p$, which maps the expectation value of a Pauli observable $\sigma$ under a quantum state $\rho$ to
$\text{Tr}\left(\sigma\mathcal{N}_{p}[\rho]\right)=(1-p)\text{Tr}\left(\sigma\rho\right).$ (10)
If the parameter $p$ would be known, the results can be corrected by dividing the noisy results by $(1-p)$
$\text{Tr}\left(\sigma\rho\right)_{\text{corrected}}=\frac{1}{(1-p)}\text{Tr}\left(\sigma\rho\right)_{\text{noisy}}.$ (11)
The main idea is thus to estimate $(1-p)$ by running a circuit with a known expectation value. This becomes particularly simple when $h=0$ or $th(t)=\pi$. Hence, in both cases, the action of the field commutes with the interaction part, i.e. $[\mathcal{U}(t),\hat{n}]=0$ for $t=0$ or $th(t)=\pi$. This leads to
+¯$\langle\bar{+}|\tilde{\mathcal{U}}^{\dagger}(t)\hat{n}\tilde{\mathcal{U}}(t)|\bar{+}\rangle=(1-p),$ (12)
where $|\bar{+}\rangle$ is the uniform superposition state obtained by applying a layer of Hadamard gates to $|0\rangle^{\otimes N}$, and $\tilde{\mathcal{U}}(t)$ is the time evolution when setting $th(t)=\pi$. Since we want the circuit to be as close as possible to the original one, we choose the field such that $h(t)=\pi/t$. Therefore, the only difference with the original circuits is the rotation angles of the field, enabling us to effectively estimate $(1-p)$ by evaluating (12).
Randomized compiling
@@
Twirling can also be used to mitigate measurement errors van den Berg et al. (2022) in the computation of expectation values. This process is performed by randomly flipping the state of the qubit before measurement, and reversing the value of the measured bit, if needed. Effectively, this process approximately diagonalize the readout transfer-matrix, making it easier to invert using the following technique, described below.
Readout error mitigation
We mitigate readout errors by calibrating the device. Since a full calibration is exponentially expensive, we adopt a sparse strategy Nation et al. (2021); Kiss et al. (2022); Yang et al. (2022) by only measuring the $N$-qubit state $|0\rangle^{\otimes N}$ and $|1\rangle^{\otimes N}$, and build the confusion matrices
@@
As stated above, measurement twirling is performed to approximately diagonalize the transfer-matrix, thus making it easier to invert.
Figure 2: Quantum simulation on IQM Garnet ($N=19$). The first two cumulants of the kink density are shown as a function of the quench time, as well as the number of Trotter steps. The blue squares denote statevector simulation, the orange points the raw data, and the red the results obtained after the full pipeline of error mitigation. The complexity of the circuits is increased by one additional Trotter step per data point. The black continuous line shows the exact evolution, while the decay rate appears as a dotted line. The inset shows the renormalization factor (purple) as a function of the circuit’s depth, resulting from the expression in (12).
III Quantum simulation
In this section, we report numerical experiments, both with quantum hardware (IQM and IBM), statevector simulation (qsim Team and Collaborators (2021)) and tensor networks (itensor), of the correlation’s decay across the quantum phase transition.
III.1 Experiments on quantum hardware
We begin by performing the quench on the IQM Garnet superconducting quantum computer Abdurakhimov et al. (2024), which consists of twenty transmon qubits and achieves a two-qubit gate fidelity of 99.5% Marxer et al. (2023). Due to the chip’s connectivity, displayed in Fig. 1, we work with a system of $N=19$ spins. We implement the second-order PF method, varying the number of Trotter steps $r$ up to 20 across different quench times $\tau_{Q}\in[0.1,20]$. The first two cumulants are presented in Fig. 2, with statevector simulations shown as blue squares, raw data as purple dots, and error-mitigated results as red dots. The continuous line represents a near-exact evolution, obtained on a statevector simulator using 2000 Trotter steps. While this is not an exact result, we choose the size of the Trotter steps such that the error is at most of the order $10^{-6}$. Higher-order moments are not displayed, as they are several orders of magnitude smaller and would require significantly more measurements to be accurately estimated.
Figure 3: Quantum simulation on IBM Kingston ($N=50$) and IBM Aachen ($N=100)$. The first cumulant of the kink density is shown as a function of the quench time, as well as the number of Trotter steps. The blue squares denote MPS simulation of the exact same circuits, the orange points the raw data, and the red the results obtained after the full pipeline of error mitigation. The complexity of the circuits is increased as one additional Trotter step per data point. The black continuous line shows a proxy for the exact evolution, which is obtained with a maximal of 100 Trotter steps, while the decay rate appears as a dotted line. The inset in the $N=50$ plot refers to the renormalization factor computed according to the expression given in (15).
We perform a second 50 spins and 100 spins experiment on the ibm__\__kingston and ibm__\__aachen Heron r2 devices. They are heavy-hexagonal lattice chips consisting of 156 qubits, with two-qubit gates fidelity between $[99.83\%,99.94\%]$. The results for the first cumulant as shown in Fig. 3, and are compared to equivalent MPS simulations. The qubits are selected such as to minimize the worst two-qubit gate fidelity. The second order cumulant can not be resolved because of the limited amounts of samples collected, which should grows as $\mathcal{O}(N^{4})$.
Error mitigation is applied using noise renormalization techniques Urbanek et al. (2021); A Rahman et al. (2022); Farrell et al. (2024); Kiss et al. (2025a); Kavaki and Lewis (2024), along with 50 instances each of Pauli Cai and Benjamin (2019) and readout van den Berg et al. (2022) twirling, and matrix-free measurement mitigation Nation et al. (2021), as described in the previous Section II.3. We observe that the renormalization factor decays exponentially fast with the depth, as shown in the inset of the lower plot of Fig. 2. Therefore, this process can not be applied for an arbitrary-depth circuit, but in this case it enables us to extend it by a factor of four. Each circuits is run 2000 times, and the error bars correspond to a $95\%$ confidence interval computed via Bayesian data augmentation, see Appendix A for details. The accuracy limit resulting from finite statistics is $10^{-5/2}$, is depicted on the corresponding axis. For the larger experiments, we noted that it was more effective to choose the renormalization factor as the maximum over every observable
$(1-p)=\max_{0\leq i<N-1}\langle\bar{+}|\tilde{\mathcal{U}}^{\dagger}(t)X_{i}X_{i+1}\tilde{\mathcal{U}}(t)|\bar{+}\rangle,$ (15)
which appears as a solid line in the corresponding inset.
As predicted by the theory, we observe a constant defect density for nearly instantaneous quenches. However, the crossover to the adiabatic regime, expected at slower quenches, is absent. This is attributed to the limitations of current quantum hardware, which lacks the capability to achieve the circuit depth necessary to suppress Trotter errors in this regime. We fit the defect density decay rate for $\tau_{Q}=1$ to $\tau_{Q}=10$, as the calculations are reliable only within this range, and obtain a decay rate of $0.68$. This value deviates from theoretical predictions due to finite-size effects. Due to the concurrent effect of boundary effect and finite numbers of interacting particles, in Fig.4 we are able to identify three typical regimes in the defect density formation. For long quenching times, the correlation length becomes comparable to the system size, therefore inducing the system to evolve adiabatically. In this regime, the density of kinks remains nearly zero and is independent of the quenching time. On the other extreme, for very short quenching times, the evolution is fully non-adiabatic, causing the density of kinks to reach a maximum value determined by the final quench point. Between these two extremes lies the KZ scaling regime, where the density of kinks follows a power-law dependence on the quenching time. It is worth noting that by the time this paper was submitted an interesting theoretical analysis about the effects related to the role of boundary conditions, endpoints, and eventually system size has been reported in Garcia and Chepiga (2024), that it is in agreement with our findings.
III.2 Computation with MPS
To support our findings on finite-size systems, we conduct numerical experiments using matrix-product states (MPS) Perez-Garcia et al. (2007) with larger spin systems to approximate results obtained in the thermodynamic limit del Campo (2018).
We begin by calculating the decay rate $\alpha$ as a function of system size by fitting the data to a model of the form $\propto\tau_{Q}^{-\alpha}$, in line with the decay predicted by the KZM. The data used for the fit lies within the range $\tau_{Q}\in[1,\tau_{f}]$, where $\tau_{f}$ is defined such that $\langle\hat{n}(\tau_{f})\rangle=1/N$. For this analysis, we perform MPS simulations with a fixed error tolerance of $10^{-10}$, based on which the smallest coefficients (from the singular value decomposition) with a lower sum of their squares than this are discarded, and 300 Trotter steps, extracting the decay rate using $10^{5}$ shots per run. As previously mentioned, expectation values tend to be small for higher-order cumulants. Therefore, our focus lies on the first three cumulants.
We show in Fig. 4 that the decay rate for the first three cumulants approaches $1/2$, as predicted by del Campo (2018). The uncertainty on higher cumulants increases, due to statistical noise.
Finally, we reconstruct the kink probability distribution $P(\langle\hat{n}\rangle)$ from the first three moments at fixed system size $N=150$ for different quench times by using the maximum entropy method Mead and Papanicolaou (1984), which is shown in Fig. 5.
Figure 5: Distribution of the kink density. The probability distribution function $P(\langle\hat{n}\rangle)$ of the kink density is shown as a function of different quench times (colors). The simulation is performed using MPS on $N=150$ spins.
We observe that the probability distributions become narrower, and the expectation value of the kink density decreases with longer quench times, approaching the decay rate of $1/2$ observed across all cumulants, which is consistent with the analytical predictions of del Campo (2018).
IV Discussions
Simulations of the critical dynamical behavior till a few years ago were only able in suitably designed laboratories. Anyway, the advent of on-demand quantum digital processing units made testing the universality of scaling laws (and their deviations) in complex systems possible for a much broader number of researchers. We perform a quantum simulation of the dynamics of a quantum phase transition using a digital superconducting quantum computer, with a circuit’s depth of up to one hundred sequential CNOT gates, employing a variety of error mitigation techniques. Using error mitigation, we increase the number of reliable Trotter steps by a factor of four, at least for the first cumulant. Computing higher cumulants requires more statistics, as they are typically order of magnitude lower. Although the cost of error mitigation generally increases exponentially with the level of noise Quek et al. (2024); Takagi et al. (2022); Schuster et al. (2024), making it unsustainable in the long term, we argue that it remains a vital tool for utilizing quantum computers before achieving full fault-tolerance. Specifically, we contend that error mitigation bridges the gap towards the era of early fault-tolerance Katabarwa et al. (2024) and facilitates the implementation of typical algorithms Lin and Tong (2022); Kiss et al. (2025b); Blunt et al. (2023) of this era. For physics-inspired problems, the use of error-mitigated simulations on digital devices is envisaged to fill in the gaps between universal simulation on classical computer clusters and special-purpose analog quantum simulators. It is worth noting that, the latter are restricted to realizing a model one at a time and strongly rely on a specific calibration of the device. Focusing on the scaling of the defects formation, the previous works conducted on analog platforms are focusing on a ‘quasi-annealing’ regime. In such a case, the results are obtained over a time window in which thermal and noise fluctuations become relevant, and the hypothesis of closed system evolution underlying the KZM could be easily violated. We confirmed, in a platform-agnostic setup, that it is possible to engineer the dynamics of critical systems, checking the finite-size effects, paving the way to the extension for general models in which classical simulations are infeasible. Our paper constitutes a step into the study of complex systems, and models beyond the TFIM requiring a not trivial decomposition of unitary dynamics are an avenue for future simulation.
We have empirically demonstrated in a finite-size TFQIM the deviations from the KZM during rapid quenches across a continuous phase transition. These deviations lead to a breakdown in the predicted power-law scaling of defect density with $\tau_{Q}$, resulting in a plateau preceding the $\tau_{Q}^{-\alpha}$ typical scaling. Additionally, we have shown how the distribution of the density of topological defects follows a binomial distribution, nearly Gaussian when the central limit theorem requirements are met Balducci et al. (2023). This model predicts that higher-order cumulants of the defect distribution scale with the quench time in a universal manner. These findings follow the pioneer line suggested by Del Campo Gómez-Ruiz et al. (2020); Zeng et al. (2023) to test the counting statistics of topological defects and the corresponding breakdown in experimental systems. We investigated finite-size effects using MPS by analyzing the decay rate as a function of system size, as well as recovering the full probability distribution. We observed that the decay rate converges to $1/2$ for the first three cumulants, approaching the value predicted in the thermodynamic limit and in pure coherent dynamics.
As a final remark, we would like to propel the use of digital quantum hardware as simulators of emergent phenomena. We have shown that even limited devices can be successfully adapted to work as versatile quantum simulators, without the requirement of having an ad hoc laboratory. We do surmise that our work empirically supports the utility and advantage (for the physics community at large) of devices of up to thousands of physical qubits, so far from the fault-tolerance and their ambitious tasks. Our work goes in the direction of testing the scaling laws in quantum critical phenomena and has broad relevance in nonequilibrium statistical mechanics, numerically proving KZM, and the breakdown of adiabatic dynamics, for the statistics of the kink-density in a TFQIM. We surmise our findings are of interest to better understand the validity of quantum simulation and quantum annealing Miessen et al. (2024), and the study of critical and thermalization phenomena in non-dissipative systems Polkovnikov et al. (2011); Andersen et al. (2025).
Acknowledgements.
The authors thank IQM and IBM for providing access to their devices. O.K. and M.G. are supported by CERN through the CERN Quantum Technology Initiative. A.M. acknowledges that this work is partially carried out under IRA Programme, project no. FENG.02.01-IP.05-0006/23, financed by the FENG program 2021-2027, Priority FENG.02, Measure FENG.02.01., with the support of the FNP. A.M. thanks the CERN Theoretical Physics Department for the hospitality.
Author contributions
A.M. designed the study. O.K. carried out the quantum simulation, while D.T. executed the tensor networks calculations. O.K. and A.M. wrote the manuscript with inputs from all authors. M.G. and A.M. supervised and coordinated the project.
Appendix A Propagation of statistical uncertainties
As described in Ref. (Hall et al., 2021, App. B), we use a Bayesian strategy to estimate the statistical uncertainties on the mitigated results. Instead of propagating the uncertainties through the whole process, we use Bayes theorem to generate an arbitrary number of experiments, compatible with the bare data, and compute their variance. The procedure is best described considering a single qubit, whose probability of obtaining $m$ measurement of the $|1\rangle$ state out of a total of $M$ trials is given by a binominal distribution
@@ Appendix A Propagation of statistical uncertainties
Here, $\alpha$, $\beta>0$ parameterize the beta distribution and represent the number of times the basis vector $i$ is measured. Note that we only sample from binary string appearing at least once in the data, to avoid sampling from exponentially many terms. After this inference phase, the following procedure can be used to determine the expected values: (i) sample a value $p_{k}^{\prime}$ from the posterior $P(p_{k}^{\prime}|m_{i})$. (ii) sample $L$ new measurements from the likelihood $P_{b}(m_{k}^{\prime};p_{k}^{\prime})$. (iii) Compute expectations values by averaging over the generate measurements. To generalized to multiple qubits, the Dirichlet distribution can be used, which is a closed-form prior for a multinominal distribution.